A SOLID PIECE OF IRON IS IN SHAPE OF CUBOID OF DIMENSIONS (49CM*33CM*24CM) IS MELTED TO FORM A SOLID SPHERE. FIND THE RADIUS OF SPHERE
Answer by mananth(16946)

(49CM*33CM*24CM)
V cuboid = L*W*H
=49*33*24
=38808 cm^3
that is the volume of the sphere
38808 = (4/3) pi * r^3
r^3= (3*38808)/(4 * pi)
r^3= 9264.72
r=21cm